OpenStudy (anonymous):

Subtract. (-4 – 2x^5 + 9x^7 – 8x^6) – (x^6 + 4x^5 + 4 + 3x^7)

OpenStudy (anonymous):

Basically, just combine the similar powers of x The two constant terms (-4) - 4 = -8 the x terms....there aren't any the x^2 terms...none of them either ... the x^5 terms (-2x^4) - 4x^5 = -6x^5 the x^6 terms (-8x^6) - x^6 = -9x^6 the x^7 terms 9x^7 - 3x^7 = 6x^7 Then just sum up the results: 6x^7 - 9x^6 - 6x^5 -8